(x/2) + 27 + x = 90 how to i solve this problem so i can be able to answer my other ones
Found 3 solutions by richwmiller, nyc_function, stanbon:
Answer by richwmiller(17219)

Just curious.
Had you looked for similar problems before posting?
Had you looked for lessons before posting?

Answer by nyc_function(2741)

I just e-mailed the solution. Did you get it?

Answer by stanbon(75887)

(x/2) + 27 + x = 90
---
Multiply thru by 2 to get rid of the denominator:
x + 54 + 2x = 180
3x = 126
x = 42
-------------------
Another way to do it:
Add (x/2)+x to get (3/2)x
Then:
(3/2)x + 27 = 90
(3/2)x = 63
(1/2)x = 21
x = 42
================
Cheers,
Stan H.